You are an equities analyst trying to value the equity of the Australian telecoms company Telstra, with ticker TLS. In Australia, listed companies like Telstra tend to pay dividends every 6 months. The payment around August is called the final dividend and the payment around February is called the interim dividend. Both occur annually. Today is mid-March 2015. TLS's last interim dividend of $0.15 was one month ago in mid-February 2015. TLS's last final dividend of $0.15 was seven months ago in mid-August 2014. Judging by TLS's dividend history and prospects, you estimate that the nominal dividend growth rate will be 1% pa. Assume that TLS's total nominal cost of equity is 6% pa. The dividends are nominal cash flows and the inflation rate is 2.5% pa. All rates are quoted as nominal effective annual rates. Assume that each month is exactly one twelfth (1/12) of a year, so you can ignore the number of days in each month. Calculate the current TLS share price.
Answer #1

Dividend Payment Times: Mid-Feb and Mid-Aug, Growth Rate = g = 1% and Cost of Equity = 6 %

Mid-Aug 2014 Dividend = $ 0.15 and Mid-Feb 2015 Dividend = $ 0.15

Present Value of the Mid-Aug Perpetual Dividends at the end of Mid-Aug 2014 = [0.15 x (1.01) / (0.06-0.01)] = $ 3.03

Value of these Dividends at Current Time (Mid-March 2015) = PV1 = 3.03 x (1.06)^(7/12) = $ 3.13476

Present Value of the Mid-Feb Perpetual Dividends at the end of Mid-Feb 2015 = [(0.15 x 1.01)/(0.06-0.01)] = $ 3.03

Value of these Dividends at Current Time (Mid-March 2015) = PV2 = 3.03 x (1.06)^(1/12) = $ 3.04475

Current Price = 3.13476 + 3.04475 = $ 6.17951 ~ $ 6.18